Overhead Variances, Four-Variance Analysis

Oerstman, Inc., uses a standard costing system and develops its overhead rates from the current annual budget. The budget is based on an expected annual output of 124,000 units requiring 496,000 direct labor hours. (Practical capacity is 516,000 hours.) Annual budgeted overhead costs total $828,320, of which $590,240 is fixed overhead. A total of 119,200 units using 494,000 direct labor hours were produced during the year. Actual variable overhead costs for the year were $261,300, and actual fixed overhead costs were $556,150.

Required:

1. Compute the fixed overhead spending and volume variances.

Fixed Overhead Spending Variance $34090 Favorable
Fixed Overhead Volume Variance $22848 Unfavorable

2. Compute the variable overhead spending and efficiency variances. Do not round intermediate calculations

Variable Overhead Spending Variance $ Unfavorable
Variable Overhead Efficiency Variance $ Unfavorable
Answer #1

1. fiexd spending variance

=actual fixed overhead -budgeted fixed overhead

=556150-590240

=34090$ favorable

fixed overhead variance

(Actual hours- standard hours)*fixed oh application rate

=[494000-(496000/124000)*119200)]*590240/496000

=17200*1.19

=20468 unfavorable

Variable Overhead Spending Variance =

Actual hours(Actual rate - standard rate)

494000[ (261300/494000) - (238080/496000)]

=24180$ unfavorable[ As the actual cost is more than budgeted the variance is unfavorable]

Variable Overhead Efficiency Variance

standard hours *standard variable overhead per hour - actual hours*standard variable overhead per hour

standard hour per unit = 496000/124000

=4 hour per unit

standard hour for actual output = 4*119200units

=476800 hours

Variable Overhead Efficiency Variance

=476800* (238080/496000) - 494000*(238080/496000)

=228864-237120

=8256$ unfavorable